OpenStudy (darkigloo):

How do I solve for x? f(x) = 7 - 15x + 9x^2 - x^3 For the answer it says : -x^2+8x-7=0 (-x+7)(x-1)=0 x=7 x=1 But how did they get to that?

OpenStudy (the_fizicx99):

You want to find the zero's? (x intercepts)

OpenStudy (darkigloo):

yes

OpenStudy (the_fizicx99):

Simply set f(x) for 0 and factor it, remember to put it in standard form though :> f(x) = -x^3 + 9x^2 - 15x + 7 0 = -x^3 + 9x^2 - 15x + 7 Now try to factor it :>

OpenStudy (darkigloo):

that's what i don't understand. how would you factor that?

OpenStudy (the_fizicx99):

Factor it by grouping :o Take the first two terms and factor that, then take the last two terms and factor those :> The neat thing is that when you factor it, you should get the same value inside the parenthesis cx

OpenStudy (darkigloo):

so -x^3+9x^2 would be -x^2(x-9)? and how would I factor -15x+7?

OpenStudy (anonymous):

you can't factor it like that, and solve it using just alg 1 for this. Do you know what the possible zeros are? Check them...

OpenStudy (the_fizicx99):

Hm, well, -1((x-1)^2(x-7)) = 0, if you multiply both sides by -1 you would get (x-1)^2(x-7) = 0 (x-1)^2 = 0 √(x-1)^2 = 0 x - 1 = 0 +1 +1 x = 1 (x-7) = 0 +7 + 7 x = 7 So you have x = {1,7}, I meant to regroup them in my last reply x.x

OpenStudy (the_fizicx99):

Here, see if this helps; http://www.sparknotes.com/math/algebra2/polynomials/section4.rhtml

OpenStudy (darkigloo):

alright..i'm trying to understand this.

OpenStudy (the_fizicx99):

Have fun cx

OpenStudy (darkigloo):

for the explanation of the answer, it says -x^2+8x-7=0 (-x+7)(x-1)=0 x=7 x=1 How did they get -x^2+8x-7=0

OpenStudy (the_fizicx99):

Hint: p/q and synthetic division

OpenStudy (darkigloo):

ohh ok. Thank you!

OpenStudy (the_fizicx99):

Yw xD